Construct a table of valued of the function using the interval of -5 to 5.

a. g(x)= x3+3x-5/x2

1
Expert's answer
2021-01-26T01:48:41-0500

Solution


  • Substitute value for x & get the corresponding "g(x)" values.
  • Interval for x values can be chosen as per the requirement. Here the interval is taken as unit 1

"\qquad\qquad
\begin{aligned}
g(-5)&=(-5)^3+3(-5)-\frac{5}{(-5)^2}\\
&=-125-15-0.2=-140.2\\
g(-4)&=(-4)^3+3(-4)-\frac{5}{(-4)^2}\\
&=-64-12-0.3125=-76.3125\\
g(-3)&=(-3)^3+3(-3)-\frac{5}{(-3)^2}\\
&=-27-9-0.5556=-36.5556\\
g(-2)&=(-2)^3+3(-2)-\frac{5}{(-2)^2}\\
&=-8-6-1.25=-15.25\\
g(-1)&=(-1)^3+3(-1)-\frac{5}{(-1)^2}\\
&=-1-3-5=-9\\
g(0)&=(0)^3+3(0)-\frac{5}{(-0)^2}\\
&=\infty\\
g(1)&=(1)^3+3(1)-\frac{5}{(1)^2}\\
&=1+3-5=-1\\
g(2)&=(2)^3+3(2)-\frac{5}{(2)^2}\\
&=8+6-1.25=12.75\\
g(3)&=(3)^3+3(3)-\frac{5}{(3)^2}\\
&=27+9-0.5556=35.4444\\
g(4)&=(4)^3+3(4)-\frac{5}{(4)^2}\\
&=64+12-0.3125=75.688\\
g(5)&=(5)^3+3(5)-\frac{5}{(5)^2}\\
&=125+15-0.2=139.8\\
\end{aligned}"

  • Then the table is